OpenStudy (anonymous):

If f(x) = 7x-4, find f(3)-2f(-2)

OpenStudy (anonymous):

where do you need help for this? what is confusing?

OpenStudy (zepp):

\(f(x)=7x-4\), find \(f(3)-2f(-2)\).

OpenStudy (zepp):

\(f(x)=7x-4\), when you are asked to find the solution of \(f(c)\) of any function, where c is a number/constant or whatever, simply plug in what you have in these parentheses in \(x\) and solve.

OpenStudy (zepp):

Now can you apply what I just told you to your problem?

OpenStudy (zepp):

@qasaali Let her try :)

OpenStudy (anonymous):

ok zepp

OpenStudy (anonymous):

i try wait

OpenStudy (zepp):

E.g.: Given \(\large f(x)=7x-4, \text{find f(c)}; f(c) = 7c-4\).

OpenStudy (anonymous):

let me help her

OpenStudy (zepp):

You can do it >:D

OpenStudy (anonymous):

i need a help guys

OpenStudy (zepp):

First can you find f(3)?

OpenStudy (zepp):

\(\large \text{Given }f(x) =7x-4, \text{Find f(c);} \\\large f(c)=7c-4\) \(\large \text{Given }f(x) =7x-4, \text{Find f(3);} \\\large \text{?}\)

OpenStudy (anonymous):

wait i almost get it

OpenStudy (anonymous):

f(3)=7(3)-4 f(3)=21-4 f(3)=17

OpenStudy (zepp):

http://puu.sh/MawP

OpenStudy (anonymous):

yeah i got 17 also

OpenStudy (zepp):

Great! but your actual problem is to find \(f(3)-2f(-2)\), now you've understood the concept, this should be easy! :D

OpenStudy (zepp):

\(f(3)-2f(-2)\) Value of f(x) when x = 3 minus Twice the value of f(x) when x = -2

OpenStudy (anonymous):

is the answer 73 ?

OpenStudy (anonymous):

it looks like 17-2(-56) = 73

OpenStudy (anonymous):

u c we have 21 for f(3) the second step would be finding 2f(-2)

OpenStudy (zepp):

\[\text{Given}~f(x) = 7x-4, \text{Find}~f(3)-2f(-2);\\\]\[\large \begin{align} &=(7(3)-4))-2(7(-2)-4))\\ &=17-2(-14-4)\\ &=17-2(-18)\\ &=17+36\\ &=53 \end{align}\]

OpenStudy (zepp):

Umm

OpenStudy (anonymous):

f(-2)=7(-2)-4 f(-2)=-14-4 f(-2)=-18*2 f(-2)=-36

OpenStudy (anonymous):

therefore=21-36= -15

OpenStudy (anonymous):

i got 53

OpenStudy (zepp):

@qasaali It's 17, not 21, and it's -(-36)

OpenStudy (zepp):

17-(-36)=17+36=53

OpenStudy (anonymous):

17-2(-18) = 53

OpenStudy (zepp):

@glares25 Correct! Good job! :)

OpenStudy (anonymous):

ok sorry for the silly mistake (-)*(-)=+ so the last answer will be 36+17=53

OpenStudy (anonymous):

THANKS A LOT GUYS :D i probably dont figured it out if u didn't help me :)

OpenStudy (zepp):

You are welcome :)